Q: a.The spot price of an investment asset that provides no income is $30 and the...

Q:

a.The spot price of an investment asset that provides no income is $30 and the risk-free rate for all maturities (with continuous compounding) is 10%. What is the two-year forward price of the asset?

b. if the forward price of the asset is 35$, how will you arbitrage?

A) Long the asset in the forward market, short sell the asset in the spot market, and invest the proceeds into a risk free bond.

B) Short the asset in the forward market, short sell the asset in the spot market, and invest the proceeds into a risk free bond.

C) Short the asset in the forward market, borrow cash, and buy the asset in the spot market.

D) Long the asset in the forward market, buy the asset in the spot market, and invest in a risk free bond.

c.  if the forward price of the asset is 40$, how will you arbitrage?

A) Long the asset in the forward market, short sell the asset in the spot market, and invest the proceeds into a risk free bond.

B) Short the asset in the forward market, short sell the asset in the spot market, and invest the proceeds into a risk free bond.

C) Short the asset in the forward market, borrow cash, and buy the asset in the spot market.

D) Long the asset in the forward market, buy the asset in the spot market, and invest in a risk free bond.
Answer #1

Price of 2 year forward should be =S0*e^(rt)=30*e^(10%*2)=36.64208274

If market price is less, buy forward otherwise sell forward

1.
A) Long the asset in the forward market, short sell the asset in the spot market, and invest the proceeds into a risk free bond.

2.
C) Short the asset in the forward market, borrow cash, and buy the asset in the spot market.